# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ Abstract
Achieving unified multi-robot coordination and motion planning in complex environments is a challenging problem. In this paper, we present a hierarchical approach to long-range coordination, which we call Stratified Topological Autonomy for Long-Range Coordination (STALC). In particular, we look at the problem of minimizing visibility to observers and maximizing safety with a multi-robot team navigating through a hazardous environment. At its core, our approach relies on the notion of a dynamic topological graph, where the edge weights vary dynamically based on the locations of the robots in the graph. To create this dynamic topological graph, we evaluate the visibility of the robot team from a discrete set of observer locations (both adversarial and friendly), and construct a topological graph whose edge weights depend on both adversary position and robot team configuration. We then impose temporal constraints on the evolution of those edge weights based on robot team state and use Mixed-Integer Programming (MIP) to generate optimal multi-robot plans through the graph. The visibility information also informs the lower layers of the autonomy stack to plan minimal visibility paths through the environment for the team of robots. Our approach presents methods to reduce the computational complexity for a team of robots that interact and coordinate across the team to accomplish a common goal. We demonstrate our approach in simulated and hardware experiments in forested and urban environments.
Index Terms:
Multi-Robot Coordination, Cooperating Robots, Multi-Robot Systems, Agent-Based Systems.
@@ I Introduction
Intelligent planning for collaborative multi-robot teams in complex environments remains a fundamental research problem. Even when restricted to discrete actions and states, the multi-robot Markov decision process can quickly become computationally intractable [1]. This situation is further exacerbated when robot teams must reason about continuous-time dynamics, dynamic obstacles, and complex environmental conditions. In this paper, we discuss our approach, Stratified Topological Autonomy for Long-Range Coordination (STALC), which seeks to develop a unified framework for coordinated motion planning with multi-robot teams. In particular, our approach investigates the problem of minimizing detection and risk when traversing through a potentially hazardous environment. To do this, we developed a hierarchical planning approach that consists of three primary levels: a high-level multi-robot graph planner, a mid-level planner for designing low-visibility paths between nodes, and a low-level planner for generating dynamically feasible, collision-free paths for each robot.
@@
Figure 1: Experimental operational scenario to minimize visibility while traversing an urban environment. A team of three Clearpath Warthogs navigate the environment while balancing the risk and rewards of moving in formations and providing overwatch for a reconnaissance mission.
Our primary contribution is a graph planning approach using Mixed-Integer Programming (MIP). Our graph planner rapidly solves for a multi-robot plan to efficiently and effectively accomplish a team goal while minimizing visibility to potential observers. We define a dynamic topological graph, which embeds the critical features of the planning problem and interactions between robots through a topological graph with dynamic edge costs. We define this graph by segmenting the environment around visibility to potential observers and between teammates. We then capture the relationships between the robots and the environment, defined in the dynamic topological graph, in our MIP formulation. In order to significantly reduce the computational complexity of this type of problem, our approach embeds multi-robot interactions directly in the edge weights instead of defining interactions through a computationally intractable number of variables for each possible state. Additionally, we detail our approach for constructing these dynamic topological graphs using environmental data and our mid-level and low-level planners for executing the multi-agent plans.
To evaluate the effectiveness of our approach, we tested our planner in a series of simulated meadow environments and on hardware in forested and urban environments. In all scenarios, we demonstrate that our planning approach can effectively navigate a team of robots through a complex environment with unified minimum-visibility planning. In an example scenario, Fig. 21 depicts a team of three Clearpath Warthog robots navigating an urban environment.
@@
This manuscript substantially extends the work in [2] which introduced a method for solving for multi-robot plans on a dynamic topological graph. In the work presented herein, we report algorithms for generating the dynamic topological graph and for utilizing the optimized plans, resulting in a full end-to-end system for multi-robot planning for coordinated maneuvers. In addition, we present hardware experiments that demonstrate our method in real-world environments.
@@ II Related Work
To address the challenges associated with multi-robot planning in complex environments, researchers have considered many different approaches to multi-robot coordination [3, 4] and cooperative multi-agent planning [1].
In many instances, researchers have addressed the problem of multi-robot coordination or task allocation. Oftentimes, solutions formulate the problem space as a set of discrete actions and states [5, 6, 7, 8, 9]. This includes conflict-based search methods [10], auction-based methods [11] and game theoretic approaches [12]. However, many of these approaches consider agents independently, ultimately ignoring interactions and cooperation between vehicles. Vehicle interactions can cause the decision space of the problem to grow exponentially, which leads many researchers to consider decomposing the problem into smaller local problems, such as in the game theoretic approach in [13]. Machine Learning (ML) approaches, often using Graph Neural Networks [14, 15], emerge as a way to solve the entire problem end-to-end while considering agent interactions [16] and coordination [17, 18]. However, ML approaches often suffer from poor generalization to environments that were not included in the training distribution and are categorically unable to provide optimality guarantees [16, 17, 18].
Mixed-Integer Programming (MIP) enables solving for end-to-end optimal solutions and has been studied in a variety of application spaces, including recharging or timed delivery [19, 20], scheduling and task allocation [21], and communication provisioning [22]. However, MIP is NP-complete, leading many approaches to require simplifying heuristics [23] and/or offline computation for optimal solutions [24].
Many researchers have also explored multi-robot motion planning in the context of obstacle avoidance or formation control. For instance, [25] uses MIP and minimum snap trajectories to design collision-free robot trajectories using a GPU. [26] generates a modified version of A to handle inter-robot constraints. In [27, 28], the authors explore methods for multi-vehicle formation control in the presence of obstacles.
Over the years, some researchers have attempted to provide a unified approach for coordination and control (e.g., [29, 30, 31, 32]). In many cases, these methods require operating on incredibly simplified and structured environments. In this paper, we propose a unified approach for multi-robot coordination and motion panning through a hierarchical planner, where each layer is informed by the information from the other layers of the hierarchy and our high-level topological graph encodes information about the underlying metric spaces. We then employ a MIP approach to generate coordinated multi-robot plans on the topological graph. In this work, we overcome many of the computational challenges with MIP-based approaches through a compact multi-agent planning formulation, and we present results rapidly solving end-to-end for optimal solutions.
The most comparable problem formulations to our approach in considering team coordination on graphs include [8] and [18]. In [8], the authors suffer greatly from the curse of dimensionality in the number of agents, reporting that they must decompose the problem into two separate problems and only consider two agents. The authors then present [18], which considers reinforcement learning for team coordination with the intention of addressing the curse of dimensionality. However, their policy for a general graph is limited to ten nodes. When considering training on only one graph, the authors report significant jumps in solution times (on the order of magnitude of minutes, which would be impractical for real-time operation) when considering graphs with 10-25 nodes and only 3-4 agents, additionally in many cases their approach cannot yield a result at all. Our graph planning formulation directly addresses the scalability to large numbers of agents. Furthermore, we present methods to formulate the problem convexly to enable solving for optimal solutions on variable graph sizes in seconds. Due to the computational complexity of this problem space, to the best of our knowledge, there are no reported works with approaches that are capable of computing solutions for the scale of problems we consider.
@@ III Problem Formulation
Our objective is to solve for a multi-robot team plan that minimizes the robot team’s visibility to observers and maximizes safety in an adversarial environment. More specifically, we aim to avoid line-of-sight observations by a static observer in such a way that we proceed tactically from one cover region to another. In addition to minimizing overall visibility, we introduce three foundational robot coordination constructs that we propose as methods to reduce the risk of traversing an adversarial environment:
  • Overwatch: One team at a protected vantage point oversees the movement of a traversing team to mitigate risk.
  • Formations: In areas of high vulnerability, moving in a formation increases awareness of the team.
  • Teaming: Moving as a team in general can provide a higher level of protection to the robots.
We relegate team coordination to planning on a high-level graph, while minimizing visibility and maintaining formation is managed by mid-range and low-level planners, respectively.

### v2

# Mixed-Integer Programming MILP Mixed-Integer Linear Programming

@@ Abstract
In this paper, we present Stratified Topological Autonomy for Long-Range Coordination (STALC), a hierarchical planning approach for coordinated multi-robot maneuvering in real-world environments with significant inter-robot spatial and temporal dependencies. At its core, STALC consists of a multi-robot graph-based planner which combines a topological graph with a novel, computationally efficient mixed-integer programming formulation to generate highly-coupled multi-robot plans in seconds. To enable autonomous planning across different spatial and temporal scales, we construct our graphs so that they capture connectivity between free-space regions and other problem-specific features, such as traversability or risk. We then use receding-horizon planners to achieve local collision avoidance and formation control. To evaluate our approach, we consider a multi-robot reconnaissance scenario where robots must autonomously coordinate to navigate through an environment while minimizing the risk of detection by observers. Through simulation-based experiments, we show that our approach is able to scale to address complex multi-robot planning scenarios. Through hardware experiments, we demonstrate our ability to generate graphs from real-world data and successfully plan across the entire hierarchy to achieve shared objectives.
@@ I Introduction
Planning for highly collaborative multi-robot teams remains a fundamental research challenge, especially when a significant amount of interaction between robot team members is required [1]. Even when restricted to discrete actions and states, such highly coupled multi-robot Markov decision processes can quickly become computationally intractable [2]. This situation is further exacerbated when robot teams must autonomously reason about continuous-time dynamics, dynamic obstacles, and complex environmental conditions. In this paper, we present Stratified Topological Autonomy for Long-Range Coordination (STALC), a unified approach for coordinated motion planning with multi-robot teams. STALC addresses challenges associated with highly-coupled multi-robot policies and planning across different temporal and spatial resolutions. In this context, autonomy refers to the ability of the robot team to operate independently, or autonomously, without human intervention.
@@
Figure 1: Experimental operational scenario to minimize visibility while traversing an urban environment. A team of three robots autonomously navigate the environment while balancing the risk and rewards of moving in formations and providing overwatch for a reconnaissance mission.
Our primary contribution is an optimal graph-based planning approach which can efficiently generate high-level multi-robot policies when there is a high-degree of interdependence between robot actions. Our approach relies on the notion of a Dynamic Topological Graph (DTG), where graph edges vary dynamically based on the state of the robot team. By ensuring that our objective function is convex when integrality constraints are relaxed, we formulate an efficient mixed-integer program that can solve for highly-coordinated multi-robot plans in seconds. Our graph-based representation of the environment affords us the opportunity to capture important features of the lower-level planning problems. To do so, we construct graphs that embed a type of problem-dependent road-map, where edges between nodes correspond to paths through free-space, and edge weights embed costs associated with those paths. We then solve the graph problem for a time-extended horizon, considering interactions over several future steps, to enable long-range coordination among robots. When executing a long-range plan, we use the underlying collision-free paths associated with our graph edges to formulate objectives for receding-horizon planning and enable more local behaviors like robot collision avoidance and formation control.
More specifically, we focus on solving a highly-coupled Multi-Robot Coordinated Reconnaissance Problem (MCRP). In this problem, the robot team must autonomously maneuver through an environment while minimizing the probability of detection by observers in expected locations in the environment. The risk of detection can be mitigated through coordinated tactics, such as providing support to team members via line-of-sight “overwatch.” To address this scenario, we formulate a visibility metric which we use to segment the environment and construct our topological graph. We test the performance of our approach to the MCRP through simulation-based experiments and demonstrate that our method achieves better computational efficiency in scaling with both graph size and team size by several orders of magnitude when compared to existing state-of-the-art approaches. In hardware experiments, we demonstrate our method’s ability to generate graphs from real-world data, generate feasible plans across the hierarchy, and realize mission objectives in real-world forested and urban environments. Fig. 1 shows a team of three robots autonomously executing a tactical plan using our approach.
@@
This manuscript substantially extends the work in [3] which introduced a method for multi-robot planning on a DTG. In the work presented herein, we report algorithms for generating the DTG and utilizing the optimized plans, resulting in a full end-to-end system for multi-robot planning for coordinated maneuvers. Our key contributions are as follows:
  • A novel, computationally-efficient mixed-integer optimization formulation capable of generating highly-coupled multi-robot plans on graphs.
  • An algorithm to construct DTGs from terrain data for the multi-robot reconnaissance problem.
  • A hierarchical planning architecture for end-to-end multi-robot planning for coordinated maneuvers.
  • An evaluation of our approach in simulation and hardware experiments and comparison to existing approaches.
@@ II Related Work
Coordination of multi-robot teams presents a major computational challenge, especially when robots must collaborate toward shared objectives across different temporal and spatial scales [4, 5]. While the coordination problem can be modeled most generally as a Multi-Robot Markov Decision Process (MMDP), these formulations scale exponentially with the number of robots [6]. To improve computational tractability, researchers have employed various abstractions, such as discretizing the state and action spaces or formulating problems that reduce the degree of robot interaction.
II-A Multi-Robot Task Allocation
Some approaches separate the multi-robot coordination problem into high-level task allocation and lower-level multi-robot motion planning problems. [1] and [7] provide comprehensive overviews of multi-robot task allocation and propose two different taxonomies. As discussed in [1], even under discrete state and action approximations, task allocation problems with significant coupling between multi-robot policies present a major combinatorial challenge. The literature for multi-robot task allocation is vast (see, [8]). In this paper, we focus primarily on approaches that reason about significant spatial interdependence among tasks and robots.
II-B Multi-Robot Coordination on Topological Graphs
Topological graphs, where nodes embed critical locations and edges represent spatial connectivity between those locations, have served as an explicit means of connecting a task-based discretization of the state space to an underlying metric space topology. Several multi-robot coordination problems assume this underlying graph structure, including the Multi-Agent Path Finding (MAPF) problem [9], the vehicle routing problem [10], and the set covering problem [11].
While some market-based [12, 13] and game-theoretic approaches have leveraged underlying graphs for spatial coordination [14], most methods for multi-robot coordination using graphs can be categorized as search-based, optimization-based, and learning-based approaches.
II-B1 Search-based Approaches
Search-based approaches to multi-robot coordination exploit problem structure or effective heuristics to search through the joint multi-robot state space. In particular, conflict-based search [15] has emerged as a powerful two-stage approach for the MAPF problem. Earlier approaches (e.g., [16]) employed A search to solve this class of problem. Heuristic search algorithms have also been employed to solve versions of the vehicle routing problem with temporal constraints (e.g., [17]). In [18], researchers present a search-based approach for solving the problem of team coordination on graphs, where robot team members can provide support to one another [19]. By proposing the notion of a joint state graph, they solve this multi-robot coordination problem via A search. In many cases, these general algorithmic approaches either rely on assumptions about problem structure (e.g., [15]), or scale poorly with the number of robots (e.g., [16, 18]).
II-B2 Optimization-based Approaches
Mixed-Integer Programming (MIP) has proved to be another effective means of generating multi-robot coordination policies on graphs. For instance, in [20], the authors present a Mixed-Integer Linear Programming (MILP) approach for coordinating heterogeneous multi-robot teams under spatial and temporal constraints. In [21], researchers use MILP to formulate a multi-robot coverage problem. MIP formulations have also proved effective for solving variants of the multi-vehicle robot routing problem with time constraints (e.g., [22]). In [23], the authors explore the problem of joint task allocation and scheduling on graphs under communication constraints. The authors of [24] and [25] use MILP formulations to solve a multi-robot non-adversarial search and MAPF problem, respectively. In [3], researchers formulate a MILP approach to solve a multi-robot coordination problem on graphs with support, similar to the problem described in [19]. While a powerful general framework, MIP is often not computationally efficient enough to be run online (e.g., [24]) and authors frequently resort to simplifying heuristics [25].
II-B3 Learning-based Approaches
Machine Learning (ML) has also been applied to address the challenges associated with multi-robot coordination on topological graphs. In particular, Graph Neural Networks [26] have emerged as a powerful means of representing multi-robot policies. In [27], the authors use GNNs to generate policies for decentralized multi-robot path planning. In [28], researchers use GNNs to address the multi-robot coverage problem, where the underlying graph is constructed from waypoints, robot positions, paths through free space, and communications constraints. Researchers have also explored more advanced GNN frameworks. For instance, in [29, 30], the authors use a type of GNN known as a Graph Attention Network (GAT) to achieve multi-robot coordination with temporal and spatial constraints.
In [31] and [32], authors use Reinforcement Learning (RL) to achieve dynamic coalition forming and routing and task allocation and scheduling, respectively. In [33], researchers use Q-learning and Proximal Policy Optimization (PPO) for multi-robot coordination on graphs, and demonstrate the ability to scale to larger numbers of robots and graph sizes when compared to the search-based method in [18].
While able to scale to larger environments and team sizes, learning-based approaches often suffer from poor generalization to environments that were not included in the training distribution and cannot provide optimality guarantees [33].
II-C Multi-Robot Task and Motion Planning
Topological graph representations of the environment can more tightly couple task allocation and motion planning. For example, [34] encodes free-space connectivity in graph edges and embeds metric path costs as edge weights, allowing task allocation to account for team travel costs. Similarly, precomputed metric plans between nodes can be used to guide local planners toward a global solution. [35] employs a hierarchy of graphs constructed using probabilistic roadmaps to solve a multi-robot routing problem. [36] presents an alternative roadmap approach to task allocation and motion planning.
Other multi-robot approaches attempt to achieve an even tighter coupling between high and low-level planning. For instance, [37] presents an approach for joint trajectory optimization goal assignment. In [38], the authors describe an approach for achieving both collision avoidance and coverage constraints. In [39], the authors leverage conflict-based search to achieve multi-robot task allocation and motion planning. Recently, RL approaches [40, 41] have demonstrated solving the joint multi-robot task and motion planning problem.
II-D Environment Segmentation and Graph Generation
An environment can be segmented into discrete regions to be represented using a topological graph. Such graph-based roadmaps have a long history for single robot planning [42] and have been constructed via a wide variety of segmentation strategies [43]. These strategies include segmenting free-space in indoor environments [44] and terrain traversability in outdoor environments [45]. In some cases, for heterogeneous robot teams, distinct graphs have been constructed to capture different robot mobility characteristics (e.g., [34]).
In this paper, we solve a joint multi-robot coordination and motion planning problem for a reconnaissance navigation task using an underlying topological roadmap. We first construct a topological graph by segmenting the environment based on visibility and traversability. We then build on the approach in [3] to efficiently solve a highly-coupled multi-robot graph-based coordination problem using MILP. During plan execution, we leverage the metric paths embedded in our graph to guide local, low-level motion planning and formation control. We show that our approach, unlike other MIP-based approaches, is computationally tractable enough to afford real-time updates, and that our roadmap-like approach can effectively guide low-level motion plans in real-world environments.
@@ III Problem Formulation
We consider a general Multi-Robot Coordinated Maneuver Problem (MCMP), where a team of $n_{A}$ homogeneous robots must coordinate over time to achieve a shared objective. For $i\in[1,n_{A}]$, the dynamics of the $i^{\textrm{th}}$ robot are given as $x_{t+1}^{i}=f(x_{t}^{i},u_{t}^{i})$, where $x_{t}^{i}\in\mathbb{R}^{n_{x}}$ are robot states and $u_{t}^{i}\in\mathbb{R}^{n_{u}}$ are robot actions at time index $t\in[1,n_{T}]$. Let the robot team state at time $t$ be $X_{t}=\{x_{t}^{1},\dots,x_{t}^{n_{A}}\}$. Given a joint state objective function $g_{x}$, a cost on individual robot actions $g_{u}$, a set of joint state constraints $c_{x}$, and input constraints $c_{u}$, we write the MCMP as follows:
$\min_{u^{i}_{t}\kern 4.09723pt\forall i,t}$ $\sum_{t=1}^{n_{T}}\bigg(g_{x}(t,X_{t},X_{D})+\sum_{i}g_{u}(u^{i}_{t})\bigg)$ (1)
s.t. $x^{i}_{t+1}=f(x^{i}_{t},u_{t}^{i}),\quad X_{1}=X_{B},$ $\forall i,t$
$c_{x}(X_{t})\leq 0,\quad c_{u}^{i}(u_{t}^{i})\leq 0,$ $\forall i,t.$
Here, $X_{D}$ is a global goal state and $X_{B}$ is the initial state of the multi-robot team. Given that $g_{x}$ is a substantially complex, environment-dependent objective function that introduces tight, long-horizon inter-robot dependencies, the optimization problem in (1) presents a fundamental combinatorial challenge and scales exponentially with $n_{A}$ [46]. As discussed in Sec. II, to improve computational tractability, the problem in (1) can be decomposed into a discrete high-level, multi-robot coordination problem and a continuous lower-level multi-robot motion planning problem.
III-A High-Level Multi-Robot Coordination
To formulate the high-level multi-robot coordination problem, we discretize the state and action spaces by introducing the notion of a Dynamic Topological Graph (DTG). Consider a weighted graph of the form $G=(V,E,W_{E})$ with a set of nodes $v\in V$, edges $e\in E$, and edge weights $w_{e}\in W_{E}$. Let the set of graph nodes, $V$, be given by the mapping $M_{v}:\mathcal{R}\mapsto V$, where $v_{j}=M_{v}(r_{j})$. $\mathcal{R}$ is a set of non-overlapping regions $r\in\mathcal{R}$, such that $r\subseteq\mathbb{R}^{2}$ and $r_{j}\cap r_{k}=\varnothing,~\forall j\neq k$.
Now let $\mathcal{T}_{\textrm{feas}}$ be the set of all feasible paths between regions, and let $\mathcal{T}_{v_{j}v_{k}}$ be a path between regions $r_{j}$ and $r_{k}$. We then define the set of directed graph edges as ordered pairs of nodes connected via free-space, $E=\{(v_{j},v_{k})\in V\times V:\exists\mathcal{T}_{v_{j}v_{k}}\in\mathcal{T}_{\textrm{feas}}\}$}.
Finally, let the state of the robot team on the graph at time $t$ be $S_{t}=\{s_{t}^{1},\dots,s_{t}^{n_{A}}\}$, where $s_{t}^{i}\in E$ is the edge occupied by the $i^{\textrm{th}}$ robot at time $t$. We can now define a DTG as follows:
Definition III.1 ( Dynamic Topological Graph (DTG)).
Given a weighted graph $G=(V,E,W_{E})$, a DTG is a weighted graph where the edge weights are defined as a function of the state $S_{t}$, such that $w_{e,t}\in W_{E}:S_{t}\rightarrow\mathbb{R}_{>0}$.
To describe the evolution of a robot state on the graph, we define the Next Edge Action Set, $A$, as follows:
Definition III.2 (Next Edge Action Set).
For a directed edge $e=(v_{j},v_{k})\in E$, the Next Edge Action Set $A$ is defined as $A(e)=\{(v_{k},v_{l})\in E:v_{l}\in V\}$.
Given Def. III.1 and III.2, we can formulate the graph-based multi-robot coordination problem as follows:
$\min_{s^{i}_{t}\kern 4.09723pt\forall i,t}$ $\sum_{t=1}^{n_{T}}\sum_{i=1}^{n_{A}}w_{s_{t}^{i},t}(S_{t})$ (2)
s.t. $s^{i}_{t+1}\in A(s^{i}_{t}),\quad\forall i,~t\in[1,n_{T}-1]$
$S_{1}=S_{B},\quad S_{n_{T}}=S_{D}.$
Here, $S_{B}$ and $S_{D}$ are the initial and final robot graph states.
III-B Local Multi-Robot Motion Planning
While this graph-based problem formulation, if solved, results in high-level coordination, we still ultimately desire to generate policies over continuous state and action spaces. Using the resulting routes from (2), we can formulate a series of reduced local optimization problems of the same form as (1). Each local optimization problem corresponds to a coalition of robots $\mathcal{C}_{e}$ sharing edge $e\in E$, where $\mathcal{C}_{e}=\{i\in\mathbb{Z}^{+}:s^{i}=e\}$ with joint state space $X^{\mathcal{C}_{e}}$. In this case $g_{x}$, $g_{u}$, $c_{x}$, and $c_{u}$ may be replaced by edge-specific functions $g_{x,e}$, $g_{u,e}$, $c_{x,e}$, and $c_{u,e}$. The goal becomes $X^{\mathcal{C}_{e}}_{G}$, the initial state becomes $X_{B}^{\mathcal{C}_{e}}$, and the time steps become $t_{e}\in[1,n_{T_{e}}]$, which correspond to the time traversing edge $e$.
III-C Multi-Robot Coordinated Reconnaissance Problem
In this paper, we focus on a particular instance of the MCMP, which we refer to as the Multi-Robot Coordinated Reconnaissance Problem (MCRP). In this scenario, a team of robots must traverse an environment to a goal location in minimum time while minimizing the risk of detection by observers at expected locations. To mitigate risk, the robot team can employ the following strategies:
  1. i)
    Minimize Visibility: Individual robots can reduce risk by minimizing visibility with respect to observers.
  2. ii)
    Provide Overwatch: Robots can oversee portions of team members’ movement through high-visibility areas.
  3. iii)
    Move in Formation: Robots can work together and move in formations through high-visibility areas.
These actions are based on foundations of tactical mission planning [47, 48, 49]. While many methods for representing risk have been explored, we employ a simple strategy of capturing risk with path costs [50].
As proposed, the MCRP is a useful exemplar of the far more general MCMP, since it introduces tight spatial coupling through strategy (ii) and encourages coalition forming in strategy (iii). We can achieve these strategies via a specific formulation of the weighted edge cost $w_{e,t}$ in a DTG. We note that, MCRP reduces to Team Coordination on Graphs with Risk Edges (TCGRE) as described in [19] when only considering strategy (ii).
Following the iTax taxonomy presented in [1], MCRP falls in the cross-schedule dependency class (XD), where robots must coordinate with each other for multi-task (MT), multi-robot tasks (MR), over a time-extended allocation (TA). In [1], this subclass is defined as XD [MT-MR-TA]. At the time of this taxonomy, the authors were not aware of any solutions to this problem. A more recent analysis of this problem in [51] additionally confirms this class of problems has not been explored further.